OpenStudy (anonymous):

I need to solve this: 6x-5z=17 5x-y+3z=-1 2x+y=-41 I know how to solve problems with all three variables in each equation, but this has me stumped. Please help!

OpenStudy (anonymous):

From the first Equation: 6x-5z=17 z=(6x-17)/5

OpenStudy (anonymous):

Substitue that in the second equation......

OpenStudy (anonymous):

And solve that equation and the third equation

OpenStudy (anonymous):

Got it?

hartnn (hartnn):

find z in terms of x from 1st equation find y in terms of x from 3rd equation put both in 2nd equation,u will be left with only x so u will get x then put in 1st equation to get z and in 3rd equation to get y

OpenStudy (anonymous):

I'm still not getting it. When I isolate either x or z in the first equation I get a decimal answer, which leads to other decimal answers. I actually have the answers I'm supposed to get (-3, -35, -7), but I don't get them whatever I do!

hartnn (hartnn):

6x-5z=17 so, 6x-17=5z so,z=(6x-17)/5 do not simplify this....put it as it is in 2nd equation....try it

OpenStudy (anonymous):

Got 23x-y+3z=-4.4.... ?

hartnn (hartnn):

how did u still get y and z u were replacing y,z in terms of x...so u must have equation only in x...

OpenStudy (anonymous):

Dang. Doing it over...

OpenStudy (anonymous):

y=-23x+2.4 Substitute it in the third one now?

hartnn (hartnn):

how did u get that?

OpenStudy (anonymous):

5x-y+3(6x-17/5)=-1 5x-y+18x-51/15=-1 23x-y-3.4=-1 23x-y=2.4 y=-23x+2.4

hartnn (hartnn):

5 is in the denominator,not in numerator...so it won't be 15 let me do it for u form 1st equation,we have z=(6x-17)/5 from 3rd equation we have y=-41-2x when we put it in 2nd,we have 5x+2x+41+3*(6x-17)/5=-1 can u solve further for x??

hartnn (hartnn):

@InTheTardis

OpenStudy (anonymous):

Okay, I got x=-1.

hartnn (hartnn):

let me find x for u 5x+2x+41+3*(6x-17)/5=-1 so 7x+3*(6x-17)/5=-42 so 35x+3*6x-3*17=-42*5 so 35x+18x=-210+51 so 53x=-159 so x=(-159)/53 so x=-3 now tell which step u did not understand,from beginning??

OpenStudy (anonymous):

I just wasn't sure what to do with the 6x-17/5. I simplified everything on the left side as best I could; I didn't think to move the variables to the right first.

hartnn (hartnn):

so did u get all steps now? or should i explain some steps,which u didn't get??

OpenStudy (anonymous):

What are the asterisks? Multiplying?

hartnn (hartnn):

yup

OpenStudy (anonymous):

Why did you multiply everything by 5?

hartnn (hartnn):

7x+3*(6x-17)/5=-42  so that i get rid of the denominator.... by multiplying 5 throughout...